# Beyond Diagonal RIS: Passive Maximum Ratio Transmission and Interference Nulling Enabler

### v1

@@ III-A Passive MRT
To maximize the desired signal power at all users simultaneously, the diagonal elements of $\mathbf{E}$ need to be considered, i.e., $\left[\mathbf{e}_{k}\right]_{k}$. Following the approach in [24], it is noted that $\boldsymbol{\omega}_{k}$ and $\exp{\left(j\theta_{k}\right)}\boldsymbol{\omega}_{k}$ for any common phase rotation $\theta_{k}\in\mathbb{R}$ lead to the same values of $|[\mathbf{e}_{k}]_{k}|$. Hence, the phase ambiguity can be exploited to make the inner product $\mathbf{h}_{k}^{T}\boldsymbol{\omega}_{k}$ real and positive without loss of optimality. This motivates us to maximize the real part of $\mathbf{h}_{k}^{T}\omega_{k}$ instead of its absolute value. Therefore, to capture the desired signal at all users, the trace of the real-valued equivalent channel will be used while relaxing the challenging unitary and symmetric constraints. Consequently, we can find a relaxed $\boldsymbol{\Theta}$ by solving the following problem
@@
The objective function can be written as $f\left(\boldsymbol{\Theta}\right)=\mathrm{Tr}\left(\mathrm{Re}\left[\mathbf{E}\right]\right)=\frac{1}{2}\left(\mathrm{Tr}\left(\mathbf{E}\right)+\mathrm{Tr}\left(\mathbf{E}^{\ast}\right)\right)=\frac{1}{2}\left(\mathrm{Tr}\left(\mathbf{E}\right)+\mathrm{Tr}\left(\mathbf{E}\right)^{\ast}\right)=\mathrm{Re}\left[\mathrm{Tr}\left(\mathbf{E}\right)\right]$. Hence, $f\left(\boldsymbol{\Theta}\right)=\mathrm{Tr}\left(\mathrm{Re}\left[\mathbf{H}\boldsymbol{\Theta}\mathbf{W}\right]\right)=\mathrm{Re}\left[\mathrm{Tr}\left(\mathbf{H}\boldsymbol{\Theta}\mathbf{W}\right)\right]$. Using the cyclic property of the trace, we can write the objective function in (6a) as $f\left(\boldsymbol{\Theta}\right)=\mathrm{Re}\left[\mathrm{Tr}\left(\mathbf{G}\boldsymbol{\Theta}\right)\right]$, where $\mathbf{G}=\mathbf{WH}\in\mathbb{C}^{N\times N}$ is the cascaded BS-BD-RIS-$U_{k}$ channel matrix. By expanding the trace and the Frobenius norm, the optimization problem can be written as
$(P2a):$ $\,\max_{\boldsymbol{\Theta}}f\left(\boldsymbol{\Theta}\right)=\mathrm{Re}\left[\sum_{i=1}^{N}\sum_{j=1}^{N}\left[\mathbf{G}\right]_{i,j}\left[\boldsymbol{\Theta}\right]_{j,i}\right]$ (7a)
@@
The objective function in (7a) can be maximized by co-phasing the terms $\left[\mathbf{G}\right]_{i,j}$ and $\left[\boldsymbol{\Theta}\right]_{j,i}$ before summation. Such problem has a closed-form solution that follows the MRT precoding [25] and the pseudo match-and-forward precoding for non-regenerative multiple-input-multiple-output (MIMO) relays [26]. Hence, the solution of $(P2a)$ is expressed as
$\left[\boldsymbol{\Theta}\right]_{j,i}={\left[\mathbf{G}\right]_{i,j}^{\ast}}\sqrt{\frac{N}{\mathrm{Tr}\left(\mathbf{GG}^{H}\right)}}$ (8)
where the normalization factor is used to satisfy the constraint in (6b). Therefore, in a matrix form, the low-complexity $\boldsymbol{\Theta}$ can be expressed as
$\boldsymbol{\Theta}=\mathbf{G}^{H}\sqrt{\frac{N}{\mathrm{Tr}\left(\mathbf{G}\mathbf{G}^{H}\right)}}$ (9)
Since (9) does not belong to the feasible set and violates the lossless and reciprocal BD-RIS constraints, i.e., $\boldsymbol{\Theta}\boldsymbol{\Theta}^{H}=\mathbf{I}$ and $\boldsymbol{\Theta}=\boldsymbol{\Theta}^{T}$; we formulate an optimization problem to find $\boldsymbol{\Theta}$ which belongs to the feasible set by minimizing the squared Frobenius distance between the low-complexity $\boldsymbol{\Theta}$ and the desired $\boldsymbol{\Theta}$. Hence,
$(P2b):$ $\,\boldsymbol{\Theta}^{\star}=\arg\min_{\boldsymbol{\Theta}}\left\|\boldsymbol{\Theta}-\mathbf{G}^{H}\sqrt{\frac{N}{\mathrm{Tr}\left(\mathbf{G}\mathbf{G}^{H}\right)}}\right\|^{2}_{F}$ (10a)
@@
Such approach is known as a matrix approximation and is used due to its ease of computation. In addition, the squared Frobenius norm is used specifically because it is reported that it maximizes the entropy for quantized channels in MIMO systems [27]. This problem can be solved efficiently following [17, Eq. (10)-(12)], which projects the low-complexity solution to the feasible set, i.e., (9) needs symmetric unitary projection. The symmetric projection of $\boldsymbol{\Theta}$ is given as [17, Eq. (10)],
$\Pi_{\mathcal{S}_{\mathrm{recip}}}\left(\boldsymbol{\Theta}\right)=0.5\left(\boldsymbol{\Theta}+\boldsymbol{\Theta}^{T}\right)$ (11)
The unitary projection of $\boldsymbol{\Theta}$ is given as [17, Eq. (11)],
$\Pi_{\mathcal{S}_{\mathrm{loss}}}\left(\boldsymbol{\Theta}\right)=\mathbf{UV}^{H}$ (12)
where its singular value decomposition (SVD) can be written as $\boldsymbol{\Theta}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{H}$, $\mathbf{U}$ and $\mathbf{V}$ are the unitary left and right singular vectors and $\boldsymbol{\Sigma}$ contains the singular values ordered in a descending order. Finally, the unitary projection of $\Pi_{\mathcal{S}_{\mathrm{recip}}}\left(\boldsymbol{\Theta}\right)$ is given as [17, Eq. (12)],
$\Pi_{\mathcal{S}_{\mathrm{loss}}}\left(\Pi_{\mathcal{S}_{\mathrm{recip}}}\left(\boldsymbol{\Theta}\right)\right)=\widehat{\mathbf{U}}\mathbf{V}^{H}$ (13)
where $\Pi_{\mathcal{S}_{\mathrm{recip}}}\left(\boldsymbol{\Theta}\right)$ is a rank deficient matrix with rank $R$ and its unitary left and right singular vectors are $\mathbf{U}=\left[\mathbf{U}_{R},\mathbf{U}_{N-R}\right]$ and $\mathbf{V}=\left[\mathbf{V}_{R},\mathbf{V}_{N-R}\right]$. Also, $\widehat{\mathbf{U}}=\left[\mathbf{U}_{R},\mathbf{V}^{\ast}_{N-R}\right]$. Finally, we can express the optimal scattering matrix in closed-form as
@@
$\boldsymbol{\Theta}^{\star}=\Pi_{\mathcal{S}_{\mathrm{loss}}}\left(\Pi_{\mathcal{S}_{\mathrm{recip}}}\left(\mathbf{G}^{H}\sqrt{\frac{N}{\mathrm{Tr}\left(\mathbf{G}\mathbf{G}^{H}\right)}}\right)\right)$ (14)
It should be noted that the signal in general will be degraded due to the symmetric unitary projection.
@@
It will be seen in the Numerical Results and Discussions Section that the symmetric unitary projection not only will degrade the signal power but also will reduce the interference power in (5).
@@
Computational Complexity: The computational complexity of (14) depends on the complexity of (9), $\mathcal{O}\left(KN^{2}+N^{2}\right)$, and is dominated by the SVD computation for (12), $\mathcal{O}\left(N^{3}\right)$. Hence, the overall complexity can be written as $\mathcal{O}\left(N^{3}+KN^{2}+N^{2}\right)=\mathcal{O}\left(N^{3}\right)$.
@@
Channel State Information Requirements: Furthermore, the cascaded BS-BD-RIS-$U_{k}$ channel matrix, $\mathbf{G}$, is needed to compute $\boldsymbol{\Theta}^{\star}$. This can be found considering the single-connected architecture by selecting the inter-impedance elements of the BD-RIS to have an infinitely large reactance. Hence, $\mathbf{G}$ can be estimated by setting $\boldsymbol{\Theta}=\mathbf{I}$ [28] and the channel estimation is simplified without separately estimating $\mathbf{H}$ and $\mathbf{W}$.
@@ III-B Passive Interference Nulling
To nullify the interference terms in (5), the equivalent channel should result in a diagonal matrix, i.e., $\mathbf{E}=\mathbf{H}\boldsymbol{\Theta}\mathbf{W}=\mathrm{diag}\left(\boldsymbol{\Lambda}\right)$ and $\boldsymbol{\Lambda}\in\mathbb{C}^{K\times 1}$ is the interference-free equivalent channels at the users. An intuitive approach to design a low-complexity $\boldsymbol{\Theta}$ that nulls the interference follows the ZF precoding. Hence,
@@
where $\mathbf{G}^{{\dagger}}=\mathbf{H}^{H}\left(\mathbf{H}\mathbf{H}^{H}\right)^{-1}\left(\mathbf{W}^{H}\mathbf{W}\right)^{-1}\mathbf{W}^{H}$ and the normalization factor is to satisfy the constraint in (6b). Nonetheless, (15) does not belong to the feasible set. Hence, the optimal unitary symmetric scattering matrix that is the closest to (15) in terms of Frobenius norm can be written in closed-form as $\boldsymbol{\Theta}^{\star}=\Pi_{\mathcal{S}_{\mathrm{loss}}}\left(\Pi_{\mathcal{S}_{\mathrm{recip}}}\left(\mathbf{G}^{{\dagger}}\sqrt{\frac{N}{\mathrm{Tr}\left(\mathbf{G}^{{\dagger}H}\mathbf{G}^{{\dagger}}\right)}}\right)\right)$.
@@
To effectively nullify the interference, we adopt an alternative approach as detailed below. The adopted approach is inspired by the interference nulling presented for D-RIS in [10]. Hence, to null the interference in a BD-RIS-enabled MU-MISO system, the equivalent channel, i.e., $\mathbf{E}$, needs to be classified into a desired/interference channels. The case is different for BD-RIS as the impinging wave on a certain RE can be also reflected by other REs due to the fully-connected nature of the BD-RIS. Therefore, we can re-write the equivalent channel as $\mathrm{vec}\left(\mathbf{E}\right)=\mathrm{vec}\left(\mathbf{H}\boldsymbol{\Theta}\mathbf{W}\right)=\mathbf{A}\mathrm{vec}\left(\boldsymbol{\Theta}\right)$, where $\mathbf{A}=\mathbf{W}^{T}\otimes\mathbf{H}\in\mathbb{C}^{K^{2}\times N^{2}}$ denotes the desired and interference channels matrix. For instance, $\mathbf{a}^{T}_{i}=[\mathbf{A}]_{i,:}$ represents the desired/interference channels for a specific user. We define $\tilde{\mathbf{a}}_{\grave{k},k}=\mathbf{a}_{i}$, for $i=(k-1)\times K+\grave{k}$. If $\grave{k}=k$, then $\tilde{\mathbf{a}}_{\grave{k},k}$ is the desired channel for $U_{k}$. Otherwise, $\tilde{\mathbf{a}}_{\grave{k},k}$ is the interference channel for $U_{k}$ from antenna $\grave{k}$. Furthermore, the interference channels matrix can be defined as
$\bar{\mathbf{A}}$ $=\left[\bar{\mathbf{A}}_{1},\bar{\mathbf{A}}_{2},\ldots,\bar{\mathbf{A}}_{K}\right]\in\mathbb{C}^{N^{2}\times K(K-1)}$ (18a)
@@
$\bar{\mathbf{A}}_{k}$ $=[\tilde{\mathbf{a}}_{1,k},\cdots,\tilde{\mathbf{a}}_{k-1,k},\tilde{\mathbf{a}}_{k+1,k},\cdots,\tilde{\mathbf{a}}_{K,k}],$
@@
In addition, the desired channels matrix is defined as $\check{\mathbf{A}}=\left[\tilde{\mathbf{a}}_{1,1},\tilde{\mathbf{a}}_{2,2},\ldots,\tilde{\mathbf{a}}_{K,K}\right]\in\mathbb{C}^{N^{2}\times K}$. To achieve $\mathbf{E}=\mathrm{diag}(\boldsymbol{\Lambda})\Leftrightarrow\mathbf{A}\mathrm{vec}(\boldsymbol{\Theta})=\mathrm{vec}(\mathrm{diag}(\boldsymbol{\Lambda}))$, the $K(K-1)$ interference channels needs to be nulled by the BD-RIS. This can be achieved by satisfying the nulling condition, which is expressed as
$\bar{\mathbf{A}}^{T}\mathrm{vec}\left(\boldsymbol{\Theta}\right)=\mathbf{0}$ (19)
In general, (19) can be seen as a system of linear equations with $K(K-1)$ complex equations and $N^{2}$ complex variables. Nonetheless, the number of effective variables is less due to the symmetry and unitary conditions on $\boldsymbol{\Theta}$, which has effectively the diagonal elements and upper/lower triangle, that is, the number of effective variables is $\frac{N}{2}(1+N)$ real variable. Therefore, (19) can be achieved if the number of real variables is greater or equal the number of real equations, i.e., $\frac{N}{2}(1+N)\geq 2K(K-1)$. Therefore, by solving the quadratic inequality and noting that $K$ and $N$ are integers, we can write $N\geq\left\lceil\frac{-1+\sqrt{16K^{2}-16K+1}}{2}\right\rceil$ which is equivalent to $N\geq 2K-1$ for $K>1$. This can be proved by showing that $\frac{-1+\sqrt{16K^{2}-16K+1}}{2}$ lies strictly between $22K-$ 2 and $12K-$ 1 for $K>1$.
Therefore, by setting $N\geq 2K-1$, $\boldsymbol{\Theta}$ that achieves (19) always exists. Such figure for $N$ is way less compared to the D-RIS which requires at least $N>2K(K-1)$ [10].
Consequently, the objective is to find $\boldsymbol{\Theta}$ that is subject to the interference nulling, BD-RIS lossless and reciprocity conditions. In other words, we have the following feasibility-check problem

### v2

@@ III-A Passive MRT
To maximize the desired signal power at all users simultaneously, the diagonal elements of $\mathbf{E}$ need to be considered, i.e., $\left[\mathbf{e}_{k}\right]_{k}$. Following the approach in [30], it is noted that $\boldsymbol{\omega}_{k}$ and $\exp{\left(j\theta_{k}\right)}\boldsymbol{\omega}_{k}$ for any common phase rotation $\theta_{k}\in\mathbb{R}$ lead to the same values of $|[\mathbf{e}_{k}]_{k}|$. Hence, the phase ambiguity can be exploited to make the inner product $\mathbf{h}_{k}^{T}\boldsymbol{\omega}_{k}$ real and positive without loss of optimality. This motivates us to maximize the real part of $\mathbf{h}_{k}^{T}\omega_{k}$ instead of its absolute value. Therefore, to capture the desired signal at all users, the trace of the real-valued equivalent channel will be used while relaxing the challenging constraints in (9c) and (9d). Consequently, we can find a relaxed $\boldsymbol{\Theta}$ by solving the following problem
@@
The objective function can be written as $f\left(\boldsymbol{\Theta}\right)=\mathrm{Tr}\left(\mathrm{Re}\left[\mathbf{E}\right]\right)=\frac{1}{2}\left(\mathrm{Tr}\left(\mathbf{E}\right)+\mathrm{Tr}\left(\mathbf{E}^{\ast}\right)\right)=\frac{1}{2}\left(\mathrm{Tr}\left(\mathbf{E}\right)+\mathrm{Tr}\left(\mathbf{E}\right)^{\ast}\right)=\mathrm{Re}\left[\mathrm{Tr}\left(\mathbf{E}\right)\right]$. Hence, $f\left(\boldsymbol{\Theta}\right)=\mathrm{Tr}\left(\mathrm{Re}\left[\mathbf{H}\boldsymbol{\Theta}\mathbf{W}\right]\right)=\mathrm{Re}\left[\mathrm{Tr}\left(\mathbf{H}\boldsymbol{\Theta}\mathbf{W}\right)\right]$. Using the cyclic property of the trace, we can write the objective function in (12a) as $f\left(\boldsymbol{\Theta}\right)=\mathrm{Re}\left[\mathrm{Tr}\left(\mathbf{G}\boldsymbol{\Theta}\right)\right]$, where $\mathbf{G}=\mathbf{WH}\in\mathbb{C}^{N\times N}$ is the cascaded BS-BD-RIS-$U_{k}$ channel matrix. In the following, we outline the passive MRT solution for the single-connected and group-connected BD-RIS architectures, where the latter captures the fully-connected architecture as well.
III-A1 Solution for Single-Connected BD-RIS
While noting that the single-connected architecture has a diagonal $\boldsymbol{\Theta}$ with unit modulus elements, we can expand $(P2a)$ as
$(P2a):$ $\,\max_{\boldsymbol{\Theta}}f\left(\boldsymbol{\Theta}\right)=\mathrm{Re}\left[\sum_{i=1}^{N}\left[\mathbf{G}\right]_{i,i}\left[\boldsymbol{\Theta}\right]_{i,i}\right]$ (13a)
@@ III-A1 Solution for Single-Connected BD-RIS
Therefore, the optimal scattering matrix that belongs to the feasible set, i.e., $\mathcal{S}_{\mathrm{SC}_{1}}$ and $\mathcal{S}_{\mathrm{SC}_{2}}$, is given as
$\left[\boldsymbol{\Theta}^{\star}\right]_{i,j}=\left\{\begin{array}[]{lr}\Pi_{\mathcal{S}_{\mathrm{unim}}}\left(\left[\mathbf{G}\right]_{i,j}^{\ast}\right),&i=j\\ 0,&i\neq j\end{array}\right.$ (14)
where
$\Pi_{\mathcal{S}_{\mathrm{unim}}}\left(\left[\mathbf{G}\right]_{i,j}^{\ast}\right)=\frac{\left[\mathbf{G}\right]_{i,j}^{\ast}}{\left|\left[\mathbf{G}\right]_{i,j}^{\ast}\right|}$ (15)
is the projection to the unit-modulus constraint shown in (3) with the minimum Euclidean distance to $\left[\mathbf{G}\right]_{i,j}^{\ast}$ [10].
Computational Complexity: The computational complexity of (14) depends on the computational complexity of computing the diagonal elements of $\mathbf{G}^{\ast}$, $\mathcal{O}\left(KN\right)$, while the normalization has a computational complexity $\mathcal{O}\left(N\right)$. Hence, the overall complexity can be written as $\mathcal{O}\left(KN+N\right)=\mathcal{O}\left(KN\right)$.
III-A2 Solution for Group-Connected BD-RIS
While noting that the group-connected architecture has a block diagonal $\boldsymbol{\Theta}$ with unitary $\boldsymbol{\Theta}_{g}$, we define ${\mathbf{G}}_{g}\triangleq\left[\mathbf{G}\right]_{N_{g}(g-1)+1:gN_{g},N_{g}(g-1)+1:gN_{g}}\in\mathbb{C}^{N_{g}\times N_{g}}$. Hence, we can expand $(P2a)$ as
$(P2a):$ $\,\max_{\boldsymbol{\Theta}}f\left(\boldsymbol{\Theta}\right)=\mathrm{Re}\left[\sum_{g=1}^{G}\sum_{i=1}^{N_{g}}\sum_{j=1}^{N_{g}}[{\mathbf{G}}_{g}]_{i,j}\left[\boldsymbol{\Theta}_{g}\right]_{j,i}\right]$ (16a)
@@ III-A2 Solution for Group-Connected BD-RIS
Consequently, the objective function in (16a) can be maximized by co-phasing the terms $[{\mathbf{G}}_{g}]_{i,j}$ and $\left[\boldsymbol{\Theta}_{g}\right]_{j,i}$ before summation. Such problem has a closed-form solution that follows the MRT precoding [31] and the pseudo match-and-forward precoding for non-regenerative MIMO relays [32]. Consequently, the closed-form solution given as
$\boldsymbol{\Theta}$ $=\mathrm{blkdiag}\left(\boldsymbol{\Theta}_{1},\ldots,\boldsymbol{\Theta}_{G}\right)$
@@
$\boldsymbol{\Theta}_{g}$ $={\mathbf{G}}_{g}^{H}\sqrt{\frac{N_{g}}{\mathrm{Tr}\left({\mathbf{G}}_{g}{\mathbf{G}}_{g}^{H}\right)}}$ (17)
Since (17) does not belong to the feasible set, i.e., $\mathcal{S}_{\mathrm{GC}_{1}}$ and $\mathcal{S}_{\mathrm{GC}_{2}}$, we project such low-complexity $\boldsymbol{\Theta}$ to the feasible set using the SVD-based symmetric unitary projection. In other words, we formulate an optimization problem to find $\boldsymbol{\Theta}$ which belongs to the feasible set by minimizing the squared Frobenius distance between the low-complexity $\boldsymbol{\Theta}_{g}$ and the desired $\boldsymbol{\Theta}_{g}$. Hence,
$(P2b):$ $\,\boldsymbol{\Theta}_{g}^{\star}=\arg\min_{\boldsymbol{\Theta}_{g}}\left\|\boldsymbol{\Theta}_{g}-\mathbf{G}_{g}^{H}\sqrt{\frac{N_{g}}{\mathrm{Tr}\left(\mathbf{G}_{g}\mathbf{G}_{g}^{H}\right)}}\right\|^{2}_{F}$ (18a)
[…]
Such approach is known as a matrix approximation and is used due to its ease of computation. In addition, the squared Frobenius norm is used specifically because it is reported that it maximizes the entropy for quantized channels in MIMO systems [33]. This problem can be solved efficiently following [21, Eq. (10)-(12)], which projects the low-complexity solution to the feasible set, i.e., (17) needs symmetric unitary projection. A closed-form symmetric projection of $\boldsymbol{\Theta}_{g}$ is given as [21, Eq. (10)],
$\Pi_{\mathcal{S}_{\mathrm{sym}}}\left(\boldsymbol{\Theta}_{g}\right)=0.5\left(\boldsymbol{\Theta}_{g}+\boldsymbol{\Theta}_{g}^{T}\right)$ (19)
The unitary projection of $\boldsymbol{\Theta}_{g}$ is given as [21, Eq. (11)],
$\Pi_{\mathcal{S}_{\mathrm{uni}}}\left(\boldsymbol{\Theta}_{g}\right)=\mathbf{UV}^{H}$ (20)
where its SVD can be written as $\boldsymbol{\Theta}_{g}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{H}$, $\mathbf{U}$ and $\mathbf{V}$ are the unitary left and right singular vectors and $\boldsymbol{\Sigma}$ contains the singular values ordered in a descending order. Finally, the SVD-based symmetric unitary projection of $\boldsymbol{\Theta}_{g}$ is given as [21, Eq. (12)],
$\Pi_{\mathcal{S}_{\mathrm{symuni}}}\left(\boldsymbol{\Theta}_{g}\right)=\widehat{\mathbf{U}}\mathbf{V}^{H}$ (21)
where $\Pi_{\mathcal{S}_{\mathrm{sym}}}\left(\boldsymbol{\Theta}_{g}\right)$ is a rank deficient matrix with rank $R$ and its unitary left and right singular vectors are $\mathbf{U}=\left[\mathbf{U}_{R},\mathbf{U}_{N-R}\right]$ and $\mathbf{V}=\left[\mathbf{V}_{R},\mathbf{V}_{N-R}\right]$. Also, $\widehat{\mathbf{U}}=\left[\mathbf{U}_{R},\mathbf{V}^{\ast}_{N-R}\right]$. Hence, we can express the optimal scattering matrix in closed-form as
$\boldsymbol{\Theta}^{\star}$ $=\mathrm{blkdiag}\left(\boldsymbol{\Theta}^{\star}_{1},\ldots,\boldsymbol{\Theta}^{\star}_{G}\right)$ (22)
$\boldsymbol{\Theta}_{g}^{\star}$ $=\Pi_{\mathcal{S}_{\mathrm{symuni}}}\left(\mathbf{G}_{g}^{H}\sqrt{\frac{N_{g}}{\mathrm{Tr}\left({\mathbf{G}_{g}}\mathbf{G}_{g}^{H}\right)}}\right)$ (23)
@@
It should be noted that the signal in general will be degraded due to the SVD-based symmetric unitary projection. It will be seen in the Numerical Results and Discussions Section that the SVD-based symmetric unitary projection not only will degrade the signal power but also will reduce the interference power in (11).
@@
Computational Complexity: The computational complexity of (23) depends on the complexity of (17), $\mathcal{O}\left(GKN_{g}^{2}+GN_{g}^{2}\right)$, and is dominated by the $G$ SVD computations for (21), $\mathcal{O}\left(GN_{g}^{3}\right)$. Hence, the overall complexity can be written as $\mathcal{O}\left(GN_{g}^{3}+GKN_{g}^{2}+GN_{g}^{2}\right)=\mathcal{O}\left(GN_{g}^{3}\right)$.
@@
Channel State Information Requirements: Furthermore, the cascaded BS-BD-RIS-$U_{k}$ channel matrix, $\mathbf{G}$, is needed to compute $\boldsymbol{\Theta}^{\star}$ for the three architectures. This can be found by selecting the inter-impedance elements of the BD-RIS to have an infinitely large reactance. Hence, $\mathbf{G}$ can be estimated by setting $\boldsymbol{\Theta}=\mathbf{I}$ [34]. Therefore, the channel estimation is simplified without separately estimating $\mathbf{H}$ and $\mathbf{W}$.
@@ III-B Passive Interference Nulling
To nullify the interference terms in (11), the equivalent channel should result in a diagonal matrix, i.e., $\mathbf{E}=\mathbf{H}\boldsymbol{\Theta}\mathbf{W}=\mathrm{diag}\left(\boldsymbol{\lambda}\right)$ and $\boldsymbol{\lambda}\in\mathbb{C}^{K\times 1}$ is the interference-free equivalent channels at the users. Considering the most flexible architecture which is the fully-connected, an intuitive approach to design a low-complexity $\boldsymbol{\Theta}$ that nulls the interference follows the ZF precoding. Hence,
@@
where $\mathbf{G}^{{\dagger}}=\mathbf{H}^{H}\left(\mathbf{H}\mathbf{H}^{H}\right)^{-1}\left(\mathbf{W}^{H}\mathbf{W}\right)^{-1}\mathbf{W}^{H}$ and the normalization factor is to satisfy the constraint in (12b). Nonetheless, (24) does not belong to the feasible set. Hence, the optimal symmetric unitary scattering matrix that is the closest to (24) in terms of Frobenius norm can be written in closed-form as $\boldsymbol{\Theta}^{\star}=\Pi_{\mathcal{S}_{\mathrm{symuni}}}\left(\mathbf{G}^{{\dagger}}\sqrt{\frac{N}{\mathrm{Tr}\left(\mathbf{G}^{{\dagger}H}\mathbf{G}^{{\dagger}}\right)}}\right)$.
@@
To effectively nullify the interference, we adopt an alternative approach as detailed below. The adopted approach is inspired by the interference nulling presented for D-RIS in [10]. Hence, to null the interference in a BD-RIS-enabled MU-MISO system, the equivalent channel, i.e., $\mathbf{E}$, needs to be classified into a desired/interference channels. Then, $\boldsymbol{\Theta}$ shall be tuned to null all interference channels. In the following, we illustrate interference nulling for the group-connected architecture which captures the fully-connected and single-connected BD-RIS architectures as a special case.
We can re-write the equivalent channel as
$\mathrm{vec}\left(\mathbf{E}\right)=\mathbf{A}_{c}\boldsymbol{\theta}_{c}$ (27)
where $\boldsymbol{\theta}_{c}\in\mathbb{C}^{GN_{g}^{2}\times 1}$ and is given as
$\boldsymbol{\theta}_{c}=\left\{\begin{array}[]{lc}\left[[\boldsymbol{\Theta}]_{1,1},[\boldsymbol{\Theta}]_{2,2},\ldots,[\boldsymbol{\Theta}]_{N,N}\right]^{T},&c=\mathrm{SC}\\ \mathrm{vec}(\boldsymbol{\Theta}),&c=\mathrm{FC}\\ \left[\mathrm{vec}(\boldsymbol{\Theta}_{1})^{T},\ldots,\mathrm{vec}(\boldsymbol{\Theta}_{G})^{T}\right]^{T},&c=\mathrm{GC}\end{array}\right.$ (28)
$\boldsymbol{\theta}_{\mathrm{SC}}\in\mathbb{C}^{N\times 1}$, $\boldsymbol{\theta}_{\mathrm{FC}}\in\mathbb{C}^{N^{2}\times 1}$, $\boldsymbol{\theta}_{\mathrm{GC}}\in\mathbb{C}^{GN_{g}^{2}\times 1}$, $\mathbf{A}_{c}\in\mathbb{C}^{K^{2}\times GN_{g}^{2}}$ denotes the desired/interference channels and is given as
$\mathbf{A}_{c}=\left\{\begin{array}[]{lc}\mathbf{W}^{T}\odot\mathbf{H},&c=\mathrm{SC}\\ \mathbf{W}^{T}\otimes\mathbf{H},&c=\mathrm{FC}\\ \left[\mathbf{W}_{1}^{T}\otimes\mathbf{H}_{1},\ldots,\mathbf{W}_{G}^{T}\otimes\mathbf{H}_{G}\right],&c=\mathrm{GC}\end{array}\right.$ (29)
${\mathbf{W}}_{g}=\left[\mathbf{W}\right]_{N_{g}(g-1)+1:gN_{g},1:K}\in\mathbb{C}^{N_{g}\times K}$, ${\mathbf{H}}_{g}=\left[\mathbf{H}\right]_{1:K,N_{g}(g-1)+1:gN_{g}}\in\mathbb{C}^{K\times N_{g}}$, $\mathbf{A}_{\mathrm{SC}}\in\mathbb{C}^{K^{2}\times N}$, $\mathbf{A}_{\mathrm{FC}}\in\mathbb{C}^{K^{2}\times N^{2}}$, $\mathbf{A}_{\mathrm{GC}}\in\mathbb{C}^{K^{2}\times GN_{g}^{2}}$. To define the desired/interference channels, we drop the subscript defining the BD-RIS architecture for notational simplicity. For instance, $\mathbf{a}^{T}_{i}=[\mathbf{A}_{c}]_{i,:}$ represents the desired/interference channels for a specific user. We define $\tilde{\mathbf{a}}_{\grave{k},k}=\mathbf{a}_{i}$, for $i=(k-1)\times K+\grave{k}$. If $\grave{k}=k$, then $\tilde{\mathbf{a}}_{\grave{k},k}$ is the desired channel for $U_{k}$. Otherwise, $\tilde{\mathbf{a}}_{\grave{k},k}$ is the interference channel for $U_{k}$ from antenna $\grave{k}$. Consequently, the interference channels matrix can be defined as
$\bar{\mathbf{A}}_{c}$ $=\left[\bar{\mathbf{A}}_{1},\bar{\mathbf{A}}_{2},\ldots,\bar{\mathbf{A}}_{K}\right]\in\mathbb{C}^{GN_{g}^{2}\times K(K-1)}$ (30a)
@@
$\bar{\mathbf{A}}_{k}$ $=[\tilde{\mathbf{a}}_{1,k},\cdots,\tilde{\mathbf{a}}_{k-1,k},\tilde{\mathbf{a}}_{k+1,k},\ldots,\tilde{\mathbf{a}}_{K,k}],$
@@
In addition, the desired channels matrix is defined as $\check{\mathbf{A}}_{c}=\left[\tilde{\mathbf{a}}_{1,1},\tilde{\mathbf{a}}_{2,2},\ldots,\tilde{\mathbf{a}}_{K,K}\right]\in\mathbb{C}^{GN_{g}^{2}\times K}$.
To achieve $\mathbf{E}=\mathrm{diag}(\boldsymbol{\lambda})\Leftrightarrow\mathbf{A}_{c}\boldsymbol{\theta}=\boldsymbol{\lambda}$, the $K(K-1)$ interference channels needs to be nulled by the BD-RIS. This can be achieved by satisfying the nulling condition, which is expressed as
$\bar{\mathbf{A}}_{c}^{T}\boldsymbol{\theta}_{c}=\mathbf{0}$ (31)
In general, (31) can be seen as a system of linear equations with $K(K-1)$ complex equations and $GN_{g}^{2}$ complex variables. Nonetheless, the number of effective variables is less due to the constraints on $\boldsymbol{\Theta}$, i.e., (9c)–(9d), which is $\frac{N(1+N_{g})}{2}$ real variables. Therefore, given that the channel realizations are random, (31) can be achieved with high probability if the number of real variables is greater or equal the number of real equations, i.e.,
$\frac{N(1+N_{g})}{2}\geq 2K(K-1)$ (32)
which can be solved for $N$ considering $K>1$ as follows
$\left\{\begin{array}[]{lc}N\geq 2K(K-1),&c=\mathrm{SC}\\ N\geq 2K-1,&c=\mathrm{FC}\\ N\geq\left\lceil\frac{4K(K-1)}{1+N_{g}}\right\rceil,&c=\mathrm{GC}\end{array}\right.$ (33)
The expression for the fully-connected architecture can be proved by solving the quadratic inequality and noting that $K$ and $N$ are integers, and $N_{g}=N$ showing that $\frac{-1+\sqrt{16K^{2}-16K+1}}{2}$ lies strictly between $22K-$ 2 and $12K-$ 1. Furthermore, (32) can be solved for $K$ as follows
$\left\{{\begin{array}[]{lc}K\leq\left\lfloor\frac{1+\sqrt{1+2N}}{2}\right\rfloor,&c=\mathrm{SC}\\ K\leq\left\lfloor\frac{N+1}{2}\right\rfloor,&c=\mathrm{FC}\\ K\leq\left\lfloor\frac{1+\sqrt{1+N(1+N_{g})}}{2}\right\rfloor,&c=\mathrm{GC}\end{array}}\right.$ (34)
Consequently, the objective is to find $\boldsymbol{\Theta}$ that is subject to the interference nulling and BD-RIS constraints. In other words, we have the following feasibility-check problem